### v1

@@ 3.2 Subhalo mass segregation
We now turn to the spatial trends we may expect to see in observational data. Han et al. (2016) studied the spatial distribution of subhaloes in dark matter-only simulations, finding that they closely follow the dark matter particles. This implies that dynamical friction does not play an important role, at least beyond $R>0.02R_{\mathrm{200m}}$. They showed that differences in the distribution of subhaloes and dark matter particles in the literature are due not to dynamical friction but to the selection of subhaloes based on present-day subhalo mass instead of mass at accretion (see also van den Bosch et al., 2016). Our selection is based on present-day stellar mass, which is expected to be less affected by these selection biases, and any such biases will also be present in observations, for which we aim to provide a reference point.
We show the present-day mass segregation of EAGLE galaxies, expressed as the ratio $m_{\mathrm{sub}}/m_{\star}$ as a function of cluster-centric distance at fixed stellar mass, in the left panel of Figure 3. Over the range $10.05<R/R_{\mathrm{200m}}<$ 1, this ratio is well described by a power law, $m_{\mathrm{sub}}/m_{\star}\propto(R/R_{\mathrm{200m}})^{0.7}$. This is significantly shallower than the relation $m_{\mathrm{sub}}/m_{\mathrm{sub}}^{\mathrm{max}}\propto(R/R_{\mathrm{200m}})^{1.3}$ for dark matter subhaloes (where $m_{\mathrm{sub}}^{\mathrm{max}}$ is the maximum mass the subhalo has had throughout cosmic time) found by Han et al. (2016), although it is close to the slope of 0.6 in the latter relation found by van den Bosch et al. (2016) in a number of $N$-body simulations. However, in doing this comparison the strong dependence on sample selection pointed out by van den Bosch et al. (2016) should be kept in mind. We further discuss the connection between stellar mass and maximal subhalo mass in Section 4. As discussed in Section 5.3, at smaller radii the flattening of the relation is influenced by EAGLE’s mass resolution, whereby only the most massive systems are present in the simulation.
@@
Recently, large optical surveys with high image quality have allowed observational measurements of the dependence of the subhalo-to-stellar mass ratio with (projected) cluster-centric distance, i.e., subhalo mass segregation, using weak gravitational lensing. Although a detailed comparison between these results can be affected by modelling and analysis differences, overall these results do suggest mild segregation (Sifón et al., 2015, 2018; Li et al., 2016; Niemiec et al., 2017; Kumar et al., 2022; Wang et al., 2023). In order to allow a comparison with observational work, we also show the dependence with projected distance, which we refer to as $R_{xy}$, [^2] Here and throughout we use projected distances measured along the $x-y$ plane. We have verified that using the $x-z$ or $y-z$ planes produces statistically identical results, as expected. in the right panel of Figure 3. The decrease in $m_{\mathrm{sub}}/m_{\star}$ is still present but is reduced due to projection effects—to roughly a factor 2 instead of 10 between $R_{xy}=R_{\mathrm{200m}}$ and $R_{xy}=0.05R_{\mathrm{200m}}$ at fixed stellar mass. The segregation observed in EAGLE is consistent with the results of both Sifón et al. (2018) and Kumar et al. (2022), although the measurements of Kumar et al. (2022) tend to lie above the EAGLE relations, while Wang et al. (2023) found stronger segregation than seen in EAGLE. As discussed by Sifón et al. (2018), this could be due to differing definitions of subhalo mass as well as stellar mass. An additional complication in the comparison to observations is that our analysis is entirely free of interlopers, i.e., galaxies mistakenly identified as cluster members. Because of the steepness of the dependence with cluster-centric distance and because central galaxies have significantly higher subhalo-to-stellar mass ratios than satellites, this contamination will tend to increase the inferred mean mass. We leave a detailed assessment of interlopers for future work.
@@
The dependence of $\langle m_{\mathrm{sub}}/m_{\star}\mid R/R_{\mathrm{200m}}\rangle$ on stellar mass means that observationally one must ensure galaxies at different cluster-centric radii have comparable stellar mass distributions. In particular, if the mean stellar mass increases with radius by roughly half a dex or more then we would not expect to observe any segregation. However this is not the case for any of the recent weak lensing measurements, where the change is at most a factor 2, i.e., less than the bin size in Figure 3. We would therefore expect the slopes seen in observations to be comparable with those in the right panel of Figure 3. Although there is considerable point-to-point variation, there is overall agreement between observations and simulations.
The strength of segregation, on the other hand, does not depend on stellar mass (i.e., all curves in Figure 3 have the same slope). This is expected if the trend is due mostly to tidal stripping, since stellar mass is more centrally-concentrated than dark matter and is therefore the component least affected by external tidal forces (Chang et al., 2013; Smith et al., 2016).
@@ 3.3 The scatter in the relation between total and stellar masses in cluster galaxies
As shown in Figure 4, the scatter in the present-day satellite SHSMR tends to decrease slightly with stellar mass, and this behaviour shows no significant dependence on cluster mass. Considering all subhaloes, the scatter is $0.43\pm 0.01$ dex at $m_{\star}=10^{9}\,\mathrm{M}_{\odot}$, while at $m_{\star}=10^{11}\,\mathrm{M}_{\odot}$ the scatter is $0.36\pm 0.04$ dex, i.e., a decrease of $(17\pm 11)$over two orders of magnitude in stellar mass. This is statistically consistent with Niemiec et al. (2022) who find no dependence of SHSMR scatter with stellar mass. The scatter in the satellite SHSMR is larger than the central SHSMR, which also decreases slowly with stellar mass, from 0.20 dex at $m_{\star}=10^{9}\,\mathrm{M}_{\odot}$ to roughly 0.15 dex at $m_{\star}=10^{11}\,\mathrm{M}_{\odot}$.
On the other hand, there is a slight but noticeable dependence of the SHSMR scatter on cluster-centric distance (Figure 5), with a scatter approaching that of centrals at the closest separations, $R\lesssim 0.05R_{\mathrm{200m}}$ (70 kpc for a $M_{\mathrm{200m}}=10^{14}\,\mathrm{M}_{\odot}$ cluster at $z=0$). The scatter increases slowly outward from $0.20\pm 0.03$ dex at $R<0.02R_{\mathrm{200m}}$ to $0.43\pm 0.02$ dex at $R>R_{\mathrm{200m}}$, when averaging over all stellar masses. As with $\langle m_{\mathrm{sub}}/m_{\star}\rangle$, this reduction in scatter may be due in part to the disruption of subhaloes, be it artificial or physical, whereby only the most massive end of the subhalo population survives very close to the cluster centre.
@@
Figure 6: Correlation between the different lookback times listed in Table 2, in Gigayears ago (Gya). Diagonal panels show kernel density estimates of each lookback time distribution, using a Gaussian kernel with $\sigma=0.1\,\mathrm{Gyr}$, and the median marked by a blue line. Orange and blue curves and dashed vertical lines show distributions and medians for satellites in $M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$ and $10^{13}\,\mathrm{M}_{\odot}<M_{\mathrm{200m}}<5\times 10^{13}\,\mathrm{M}_{\odot}$. Each off-diagonal panel in the lower triangle shows the corresponding two-dimensional occupation; the color scale runs from white (zero) to a color given by the Pearson’s $r$ correlation between the two times, shown in the corner of each panel, and coded in the bottom colour bar. The brightness scaling is consistent across panels, with the darkest cells containing 2.5% of the population or more. Each panel in the upper triangle shows the logarithm of the mean subhalo-to-stellar mass ratio at $z=0$ in each cell, as coded in the upper colour bar. Numbers in the corners of each panel also correspond to the correlation coefficients.
@@ 4.1 Milestones in the life of a satellite galaxy
We identify the following key moments [^3] For each of these moments we use $i$ to refer to the snapshot number (increasing with cosmic time), $t$ to refer to the time, and $z$ to the redshift, depending on the context. in the life of a satellite galaxy: after being formed, we label $i_{\mathrm{cent}}$ the last snapshot in the simulation when a galaxy was labelled as central and $i_{\mathrm{sat}}$ the first snapshot it was labelled a satellite of any host, independent of the host’s mass. We also label the first time a galaxy is identified as a satellite of its current host as $i_{\mathrm{infall}}$, and the first snapshot after which it never leaves its current host again as $i_{\mathrm{acc}}$. By definition, then, $z_{\mathrm{sat}}\geq\{z_{\mathrm{cent}},z_{\mathrm{infall}}\}\geq z_{\mathrm{acc}}$. Additionally, we register the snapshots where each subhalo reached maximal total mass ($i_{\mathrm{sub,max}}$), stellar mass ($i_{\mathrm{\star,max}}$), and gas mass ($i_{\mathrm{gas,max}}$). Throughout, we use “infall” to refer to infall to the $z=0$ host specifically. The times $t_{\mathrm{sat}}$ and $t_{\mathrm{infall}}$ are sometimes referred to as first and last infall, and are analogous to $t_{\mathrm{branch}}$ and $t_{\mathrm{main}}$ in Bahé et al. (2019). [^4] Bahé et al. (2019) defined $t_{\mathrm{branch}}$ and $t_{\mathrm{main}}$ as the middle of the relevant snapshot interval, whereas we used the specified snapshot time. This introduces typically a $\approx 30$ Myr difference between those authors’ and our definitions at $z<1$ and differences much smaller at earlier times; these are insignificant for our purposes. $t_{\mathrm{infall}}$ is also similar to the residence time defined by O’Neil et al. (2023), except they defined it as the time since the subhalo first crossed $R_{\mathrm{200m}}$, while we define it based on whether it is gravitationally bound as calculated by hbt+.
Table 2: Statistics for present-day satellite galaxies’ characteristic lookback times in Gigayears ago (Gya). Brackets show redshifts corresponding to the quoted lookback times. Times are, from top to bottom: sat: first time as a satellite (see text for caveats), cent: last time as a central, infall: first time labelled as a satellite of current host, acc: definitively labelled as a satellite of current host. We also list the times of maximum stellar, subhalo (i.e., total), and gas masses.
@@
We summarize the median and 50% and 90% ranges for all characteristic times in Table 2. For instance, the time elapsed between being first labelled a satellite of its current host, $t_{\mathrm{infall}}$, and definitively so, $t_{\mathrm{acc}}$, is roughly 1 Gyr. This means we are excluding from our sample galaxies that have entered a halo but have temporarily left it until they re-enter after apocenter, so-called backsplash galaxies (Ludlow et al., 2008; Diemer, 2021; Fong & Han, 2021). We make no attempt to include these galaxies in our sample.
@@
We take a closer look at these characteristic times in Figure 6, which shows joint histograms in the bottom-left corner plot and the logarithm of the mean subhalo-to-stellar mass ratio at $z=0$ in the upper-right corner, as a function of pairs of lookback times. A few interesting features can be noted. For instance, most panels in the bottom-left triangle show that most galaxies occupy the $y\approx x$ zone—that is, most of these events happen within a few hundred Myr of other events labelled here: a present-day satellite typically spent little or no time as satellites of other host ($t_{\mathrm{acc}}\approx t_{\mathrm{infall}}\approx t_{\mathrm{cent}}\approx t_{\mathrm{sat}}$). The time of maximal subhalo mass is also highly correlated with these events, most notably with the last time a satellite was labelled a central. This is expected since, no matter the details, after this time a galaxy is or has been under the influence of a larger host and can therefore be subject to mass loss processes.
@@
There are three notable exceptions to this strong correspondence between events. The first is the time of maximal gas mass, which is $t_{\mathrm{gas,max}}^{\mathrm{lookback}}\sim 6-10\,\mathrm{Gya}$ mostly independent of other events. Like $t_{\mathrm{sub,max}}$, $t_{\mathrm{gas,max}}$ is most (though still quite poorly) correlated with $t_{\mathrm{cent}}$. The most important relation between maximal gas mass and other characteristic times is that $t_{\mathrm{gas,max}}$ is the first event to take place for the vast majority of galaxies. Given that stellar mass continues to grow well past this time, the reason for the early $t_{\mathrm{gas,max}}$ is not so much gas stripping but the termination of the gas supply. While some level of quenching is observed in low-mass host (e.g., Brown et al., 2017; Zinger et al., 2018), this is consistent with the finding in some simulations (including EAGLE) that it requires a massive host to fully shut down star formation (e.g., Pallero et al., 2022; Hough et al., 2023).
@@
The second exception is the time of maximal stellar mass, which is by and large the last event to take place. When this happens can be divided into two populations: most galaxies attained their maximal stellar mass almost 3 Gyr after their maximal subhalo mass (although this difference increases at later times); and a second population—15% of satellites—have maximal amounts of stellar mass today, i.e., $t_{\mathrm{\star,max}}^{\mathrm{lookback}}=0.0$ Gya. This population is naturally interpreted as consisting of satellites that have not yet reached their maximal stellar mass, and as shown in Figure 6 can have fallen into their present host as much as 6 Gya. The median infall time of these satellites is 2.3 Gya, compared to 6.2 Gya for the rest of the population; similarly, they were last classified as centrals 2.5 Gya compared to 7.0 Gya for the rest of the population. These “maximal stellar mass” satellites are four times as massive (in total subhalo mass) and reside in clusters 44% as massive (median values) as the rest of the population. This results from a combination of these satellites being preferentially more massive while also having spent less time under the influence of a massive host. As seen in the diagonals of Figure 6, the time of maximal stellar mass displays the largest difference between satellites in low- and high-mass clusters today: the fraction of satellites with maximal stellar masses at $z=0$ is 10% in high-mass ($M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$) clusters and 42% in low-mass ($M_{\mathrm{200m}}<5\times 10^{13}\,\mathrm{M}_{\odot}$) clusters, and the median time of maximal stellar mass occurs 4.5 Gyr earlier for satellites currently residing in massive clusters compared to those in low-mass clusters. These trends are qualitatively consistent with trends for quenched fractions found by e.g., Park et al. (2023).
@@
The third exception concerns $t_{\mathrm{sat}}$: 906 satellites—a full 32% of the present-day satellite population—became a satelllite more than 12 Gya (i.e., at most 1.3 Gyr after their formation in the simulation). Unlike the case for the satellites currently at their maximal stellar mass time, these “early satellite” population does not appear any different from the rest of the satellite population: even though their $t_{\mathrm{sat}}$ happened 5 Gyr earlier than the general population, no other characteristic time follows a significantly different distribution from the rest of the population, with the mean characteristic times being different by more than 0.1 Gyr, including $t_{\mathrm{cent}}$. Similarly, both their total and stellar masses are statistically identical to those galaxies that became satellites for the first time later on, both at the present day and at all milestones; they reside in hosts with the same median mass today; and occur in the same proportion in low- and high-mass clusters. All these early satellites were born as centrals, and while for many of them it appears the assignment as a satellite may be an algorithmic glitch—after only one or two snapshots they return to being centrals for a long time—this is not true for all of them. We therefore choose to leave $t_{\mathrm{sat}}$ as is for these galaxies. However, the peculiar distribution must be kept in mind in any discussions involving $t_{\mathrm{sat}}$.
@@
We also note that $t_{\mathrm{infall}}$ shows a higher correlation with both the times of maximal subhalo and stellar mass than $t_{\mathrm{acc}}$, which is why—coupled with the very high correlation of $r=0.86$ between $t_{\mathrm{infall}}$ and $t_{\mathrm{acc}}$—we choose to focus the later discussion on $t_{\mathrm{infall}}$ and ignore $t_{\mathrm{acc}}$ from now on. In words, once a galaxy has been identified as a satellite of its current host it starts to become influenced by it; whether at some later time hbt+ deems it to have left its host for some time (which, as discussed by Diemer, 2021, is probably a true feature of membership to a cluster by most definitions) is of secondary importance regarding the galaxy’s mass content.

### v2

@@ 3.2 Subhalo mass segregation
We now turn to the spatial trends we may expect to see in observational data. Han et al. (2016) studied the spatial distribution of subhaloes in dark matter-only simulations, finding that they closely follow the dark matter particles. This implies that dynamical friction does not play an important role, at least for low mass subhaloes or beyond $R>0.02R_{\mathrm{200m}}$. They showed that differences in the distribution of subhaloes and dark matter particles in the literature are due not to dynamical friction but to the selection of subhaloes based on present-day subhalo mass instead of mass at accretion (see also van den Bosch et al., 2016; He2023). Our selection is based on present-day stellar mass, which is expected to be less affected by these selection biases. In addition, any such biases will also be present in observations, for which we aim to provide a reference point.
We show the present-day mass segregation of EAGLE galaxies, expressed as the mean ratio $m_{\mathrm{sub}}/m_{\star}$ as a function of cluster-centric distance at fixed stellar mass, in the left panel of Figure 3. Over the range $10.05<R/R_{\mathrm{200m}}<$ 1, this ratio is well described by a power law, $m_{\mathrm{sub}}/m_{\star}\propto(R/R_{\mathrm{200m}})^{0.7}$. At smaller radii the apparent segregation diminishes as a consequence of numerical resolution, which also reduces the scatter in the SHSMR (discussed in Section 3.3, see Figure 5). At larger radii, $R>R_{\mathrm{200m}}$, we see a turnover in the SHSMR, which can be attributed to the presence of galaxies that already reached pericentre and whose apocentre lies outside $R_{\mathrm{200m}}$ (i.e. these galaxies are currently far from the cluster centre but have already suffered significant mass loss). We return to these points in Section 5.
The cluster-centric distance dependence of $\langle m_{\mathrm{sub}}/m_{\star}\rangle$ is similar to that of the ratio between present-day and maximum subhalo mass, $m_{\mathrm{sub}}^{\mathrm{max}}$, commonly measured for dark matter subhaloes in $N$-body simulations. We find $\langle m_{\mathrm{sub}}/m_{\mathrm{sub}}^{\mathrm{max}}\rangle\propto(R/R_{\mathrm{200m}})^{0.6}$ in EAGLE, and the same slope was found by van den Bosch et al. (2016) in a number of $N$-body simulations. [^2] This value is not reported explicitly by van den Bosch et al. (2016) but can be extracted from the lower-left panel of their Figure 3. This similarity between the dependence of $m_{\mathrm{sub}}/m_{\star}$ and $m_{\mathrm{sub}}^{\mathrm{max}}$ with cluster-centric distance is a consequence of the resilience of stellar mass to external forces, as discussed in more detail in later sections. This slope is different from that found by Han et al. (2016) for satellites in massive haloes (see also He2023). However, we point out that Han et al. (2016) treated selection effects carefully in order to extract the parent subhalo population while we report the slope at face value in an attempt to reproduce the observational situation. Furthermore, because we are mostly concerned with a comparison with observations, we report the mean values, $\langle m_{\mathrm{sub}}/m_{\mathrm{sub}}^{\mathrm{max}}\rangle$; instead, since the goal of Han et al. (2016) was to understand the underlying subhalo population they report the median in logarithmic space, $\mathrm{median}[\log(m_{\mathrm{sub}}/m_{\mathrm{sub}}^{\mathrm{max}})]$. Using the same statistic as Han et al. (2016) we find a slope of 0.8, closer to (but still different than) the latter authors. The remaining difference with the slope of 1.0 from Han et al. (2016) may be due in part to selection effects and in part to satellite galaxies being somewhat more resistant to tidal stripping than dark matter subhaloes because of the increased central density.
@@
Recently, large optical surveys with high image quality have allowed observational measurements of the dependence of the subhalo-to-stellar mass ratio with (projected) cluster-centric distance, i.e. subhalo mass segregation, using weak gravitational lensing. Although a detailed comparison between these results can be affected by modelling and analysis differences, overall these results do suggest mild segregation (Sifón et al., 2015, 2018; Li et al., 2016; Niemiec et al., 2017; Kumar et al., 2022; Wang et al., 2023). In order to allow a comparison with observational work, we also show the dependence with projected distance, which we refer to as $R_{xy}$, [^3] Here and throughout we use projected distances measured along the $x-y$ plane. We have verified that using the $x-z$ or $y-z$ planes produces statistically identical results, as expected. in the right panel of Figure 3. The decrease in $m_{\mathrm{sub}}/m_{\star}$ is still present but is reduced due to projection effects—to roughly a factor 2 instead of 10 between $R_{xy}=R_{\mathrm{200m}}$ and $R_{xy}=0.05R_{\mathrm{200m}}$ at fixed stellar mass. The segregation observed in EAGLE is consistent with the results of both Sifón et al. (2018) and Kumar et al. (2022), although the measurements of Kumar et al. (2022) tend to lie above the EAGLE relations, while Wang et al. (2023) found stronger segregation than seen in EAGLE. As discussed by Sifón et al. (2018), this could be due to differing definitions of subhalo mass as well as stellar mass. An additional complication in the comparison to observations is that our analysis is entirely free of interlopers, i.e. galaxies mistakenly identified as cluster members. Because of the steepness of the dependence with cluster-centric distance and because central galaxies have significantly higher subhalo-to-stellar mass ratios than satellites, this contamination will tend to increase the inferred mean mass. We leave a detailed assessment of interlopers for future work.
We note that some works have studied the segregation of mean stellar mass alone rather than the subhalo-to-stellar mass ratio. For instance, Roberts2015 found that segregation of stellar mass decreases with increasing cluster mass, and that low-mass galaxies are more strongly segregated. At cluster masses typical of our sample, they find a slope of roughly $-0.02$ dex, i.e. mean stellar mass increasing slightly towards the cluster centre. While we do not attempt to match both samples in detail, we measure an overall slope of $\log m_{\star}\propto-0.03(R/R_{\mathrm{200m}})$, consistent with Roberts2015.
@@
The strength of the segregation does not depend on stellar mass (i.e. all curves in Figure 3 have the same slope). This is expected if the trend is due mostly to tidal stripping, since stellar mass is more centrally concentrated than dark matter and is therefore the component least affected by external tidal forces (Chang et al., 2013; Smith et al., 2016), although the normalization changes following the SHSMR in Figure 2. This latter dependence means that observationally one must ensure galaxies at different cluster-centric radii have comparable stellar mass distributions. In particular, if the mean stellar mass increases with radius by roughly half a dex or more then we would not expect to observe any segregation. However this is not the case for any of the recent weak lensing measurements, where the change is at most a factor 2, i.e. less than the bin size in Figure 3. We would therefore expect the slopes seen in observations to be comparable with those in the right panel of Figure 3. Although there is considerable point-to-point variation, there is overall agreement between observations and simulations.
We also show in Figure 3 the mean $m_{\mathrm{sub}}/m_{\star}$ of central galaxies in each stellar mass bin (cf. Figure 2). Satellites have lower SHSMR than centrals at all radii. As discussed in Section 5, this is a consequence of pre-processing, i.e. the influence of lower-mass host haloes on the galaxies’ mass content before infall to the present host, which means satellites have lower masses than centrals even outside $R_{\mathrm{200m}}$.
@@ 3.3 The scatter in the relation between total and stellar masses in cluster galaxies
As shown in Figure 4, the scatter in the present-day satellite SHSMR tends to decrease slightly with stellar mass, and this behaviour shows no significant dependence on cluster mass. Considering all subhaloes, the scatter is $0.43\pm 0.01$ dex at $m_{\star}=10^{9}\,\mathrm{M}_{\odot}$, while at $m_{\star}=10^{11}\,\mathrm{M}_{\odot}$ the scatter is $0.36\pm 0.04$ dex, i.e. a decrease of $(17\pm 11)$over two orders of magnitude in stellar mass. This is statistically consistent with Niemiec et al. (2022) who find no dependence of SHSMR scatter with stellar mass. The scatter in the satellite SHSMR is larger than the central SHSMR, which also decreases slowly with stellar mass, from 0.20 dex at $m_{\star}=10^{9}\,\mathrm{M}_{\odot}$ to roughly 0.15 dex at $m_{\star}=10^{11}\,\mathrm{M}_{\odot}$.
On the other hand, there is a slight but noticeable dependence of the SHSMR scatter on cluster-centric distance (Figure 5), with a scatter approaching that of centrals at the closest separations, $R\lesssim 0.05R_{\mathrm{200m}}$ (70 kpc for a $M_{\mathrm{200m}}=10^{14}\,\mathrm{M}_{\odot}$ cluster at $z=0$). The scatter increases slowly outwards from $0.20\pm 0.03$ dex at $R<0.02R_{\mathrm{200m}}$ to $0.43\pm 0.02$ dex at $R>R_{\mathrm{200m}}$, when averaging over all stellar masses. As with $\langle m_{\mathrm{sub}}/m_{\star}\rangle$, this reduction in scatter may be due in part to the disruption of subhaloes, be it artificial or physical, whereby only the most massive end of the subhalo population survives very close to the cluster centre.
@@
Figure 6: Correlation between the different lookback times listed in Table 2 (in units of Gya). The diagonal panels show kernel density estimates of each lookback time distribution, using a Gaussian kernel with $\sigma=0.1\,\mathrm{Gyr}$, and the median marked by a blue line. The orange and blue curves and dashed vertical lines show distributions and medians for satellites in $M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$ and $10^{13}\,\mathrm{M}_{\odot}<M_{\mathrm{200m}}<5\times 10^{13}\,\mathrm{M}_{\odot}$. Each off-diagonal panel in the lower triangle shows the corresponding two-dimensional occupation; the colour scale runs from white (zero) to a colour given by the Spearman correlation coefficient between the two times, shown in the corner of each panel, and coded in the bottom colour bar. The brightness scaling is consistent across panels, with the darkest cells containing 2.5% of the population or more. Each panel in the upper triangle shows the logarithm of the mean subhalo-to-stellar mass ratio at $z=0$ in each cell, as coded in the upper colour bar. The numbers in the corners of each panel correspond to the correlation coefficients.
@@ 4.1 Milestones in the life of a satellite galaxy
We identify the following key moments [^4] For each of these moments we use $i$ to refer to the snapshot number (increasing with cosmic time), $t$ to refer to the time, and $z$ to the redshift, depending on the context. in the life of a satellite galaxy: after being formed, we label $i_{\mathrm{cent}}$ the last snapshot in the simulation when a galaxy was labelled as central and $i_{\mathrm{sat}}$ the first snapshot it was labelled a satellite of any host, independent of the host’s mass. We also label the first time a galaxy is identified as a satellite of its current host as $i_{\mathrm{infall}}$, and the first snapshot after which it never leaves its current host again as $i_{\mathrm{acc}}$. By definition, then, $z_{\mathrm{sat}}\geq\{z_{\mathrm{cent}},z_{\mathrm{infall}}\}\geq z_{\mathrm{acc}}$. Additionally, we register the snapshots where each subhalo reached maximum total mass ($i_{\mathrm{sub,max}}$), stellar mass ($i_{\mathrm{\star,max}}$), and gas mass ($i_{\mathrm{gas,max}}$). Throughout, we use “infall” to refer to infall to the $z=0$ host specifically. The times $t_{\mathrm{infall}}$ and $t_{\mathrm{acc}}$ are sometimes referred to as first and last infall, and are analogous to $t_{\mathrm{branch}}$ and $t_{\mathrm{main}}$ in Bahé et al. (2019). [^5] Bahé et al. (2019) defined $t_{\mathrm{branch}}$ and $t_{\mathrm{main}}$ as the middle of the relevant snapshot interval, whereas we used the specified snapshot time. This introduces typically a $\approx 30$ Myr difference between those authors’ and our definitions at $z<1$ and differences much smaller at earlier times; these are insignificant for our purposes. $t_{\mathrm{infall}}$ is also similar to the residence time defined by O’Neil et al. (2023), except they defined it as the time since the subhalo first crossed $R_{\mathrm{200m}}$, while we define it based on whether it is gravitationally bound as calculated by hbt+.
Table 2: Statistics for present-day satellite galaxies’ characteristic lookback times in Gigayears ago (Gya). Brackets show redshifts corresponding to the quoted lookback times. Times are, from top to bottom: sat: first time as a satellite (see text for caveats), cent: last time as a central, infall: first time labelled as a satellite of current host, acc: definitively labelled as a satellite of current host. We also list the times of maximum stellar, subhalo (i.e. total), and gas masses.
@@
We summarize the median and 50% and 90% ranges for all characteristic times in Table 2. For instance, the time elapsed between being first labelled a satellite of its current host, $t_{\mathrm{infall}}$, and definitively so, $t_{\mathrm{acc}}$, is roughly 1 Gyr. That is, some galaxies entered a halo in the recent past (typically within the last Gyr), have temporarily left it, and have not had the time to re-enter their host (once they reach apocentre) by $z=0$, so-called backsplash galaxies (Ludlow et al., 2008; Diemer, 2021; Fong & Han, 2021). Backsplash galaxies are labelled as centrals by hbt+ at $z=0$ and are therefore excluded from our sample of satellites. We make no attempt to include them.
@@
We take a closer look at these characteristic times in Figure 6, which shows joint histograms in the bottom-left corner plot and the logarithm of the mean subhalo-to-stellar mass ratio at $z=0$ in the upper-right corner, as a function of pairs of lookback times. A few interesting features can be noted. For instance, most panels in the bottom-left triangle show that most galaxies occupy the $y\approx x$ zone—that is, most of these events happen within a few hundred Myr of other events labelled here: a present-day satellite typically spent little or no time as satellites of other host ($t_{\mathrm{acc}}\approx t_{\mathrm{infall}}\approx t_{\mathrm{cent}}\approx t_{\mathrm{sat}}$). The time of maximum subhalo mass is also highly correlated with these events, most notably with the last time a satellite was labelled a central. This is expected since, no matter the details, after this time a galaxy is or has been under the influence of a larger host and can therefore be subject to mass loss processes.
@@
There are three notable exceptions to this strong correspondence between events. The first is the time of maximum gas mass, which is $t_{\mathrm{gas,max}}^{\mathrm{lookback}}\sim 6-10\,\mathrm{Gya}$ mostly independent of other events. Like $t_{\mathrm{sub,max}}$, $t_{\mathrm{gas,max}}$ is most (though still quite poorly) correlated with $t_{\mathrm{cent}}$. The most important relation between maximum gas mass and other characteristic times is that $t_{\mathrm{gas,max}}$ is the first event to take place for the vast majority of galaxies. Given that stellar mass continues to grow well past this time, the reason for the early $t_{\mathrm{gas,max}}$ is not so much gas stripping but the termination of the gas supply. While some level of quenching is observed in low-mass host (e.g. Brown et al., 2017; Zinger et al., 2018), this is consistent with the finding in some simulations (including EAGLE) that it requires a massive host to fully shut down star formation (e.g. Pallero et al., 2022; Hough et al., 2023).
@@
The second exception is the time of maximum stellar mass, which is by and large the last event to take place. When this happens can be divided into two populations: most galaxies attained their maximum stellar mass almost 3 Gyr after their maximum subhalo mass (although this difference increases at later times); and a second population—15% of satellites—have maximum amounts of stellar mass today, i.e. $t_{\mathrm{\star,max}}^{\mathrm{lookback}}=0.0$ Gya. This population is naturally interpreted as consisting of satellites that have not yet reached their maximum stellar mass, and as shown in Figure 6 can have fallen into their present host as much as 6 Gya. The median infall time of these satellites is 2.3 Gya, compared to 6.2 Gya for the rest of the population; similarly, they were last classified as centrals 2.5 Gya compared to 7.0 Gya for the rest of the population. These ‘maximum stellar mass’ satellites are four times as massive (in total subhalo mass) and reside in clusters 44% as massive (median values) as the rest of the population. This results from a combination of these satellites being preferentially more massive while also having spent less time under the influence of a massive host. As seen in the diagonals of Figure 6, the time of maximum stellar mass displays the largest difference between satellites in low- and high-mass clusters today: the fraction of satellites with maximum stellar masses at $z=0$ is 10% in high-mass ($M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$) clusters and 42% in low-mass ($M_{\mathrm{200m}}<5\times 10^{13}\,\mathrm{M}_{\odot}$) clusters, and the median time of maximum stellar mass occurs 4.5 Gyr earlier for satellites currently residing in massive clusters compared to those in low-mass clusters. These trends are qualitatively consistent with trends for quenched fractions found by e.g. Park et al. (2023).
@@
The third exception concerns $t_{\mathrm{sat}}$: 906 satellites—a full 32% of the present-day satellite population—became a satelllite more than 12 Gya (i.e. at most 1.3 Gyr after their formation in the simulation). Unlike the case for the satellites currently at their maximum stellar mass time, this ‘early satellite’ population does not appear any different from the rest of the satellite population: even though their $t_{\mathrm{sat}}$ happened 5 Gyr earlier than the general population, no other characteristic time follows a significantly different distribution from the rest of the population, with the mean characteristic times being different by more than 0.1 Gyr, including $t_{\mathrm{cent}}$. Similarly, both their total and stellar masses are statistically identical to those galaxies that became satellites for the first time later on, both at the present day and at all milestones; they reside in hosts with the same median mass today; and occur in the same proportion in low- and high-mass clusters. All these early satellites were born as centrals, and while for many of them it appears the assignment as a satellite may be an algorithmic glitch—after only one or two snapshots they return to being centrals for a long time—this is not true for all of them. We therefore choose to leave $t_{\mathrm{sat}}$ as is for these galaxies. However, its peculiar distribution must be kept in mind in any discussions involving $t_{\mathrm{sat}}$.
@@
We also note that $t_{\mathrm{infall}}$ shows a higher correlation with both the times of maximum subhalo and stellar mass than $t_{\mathrm{acc}}$, which is why—coupled with the very high correlation of $r=0.86$ between $t_{\mathrm{infall}}$ and $t_{\mathrm{acc}}$—we choose to focus the later discussion on $t_{\mathrm{infall}}$ and ignore $t_{\mathrm{acc}}$ from now on. In other words, once a galaxy has been identified as a satellite of its current host it starts to become influenced by it; whether at some later time hbt+ deems it to have left its host for some time (which, as discussed by Diemer, 2021, is probably a true feature of membership to a cluster by most definitions) is of secondary importance regarding the galaxy’s mass content.